# LANCE: Low Rank Activation Compression for Efficient On-Device Continual Learning

### v1

@@ 1 Introduction
Several approaches have been proposed to address this issue. Checkpointing (Chen et al., 2016) reduces memory usage at the cost of extra recomputation; reversible networks (Gomez et al., 2017) eliminate the need to store activations but require specialized architectures; and quantization or activation pruning (Chen et al., 2021; Barley & Fröning, 2023; Jiang et al., 2022) compress or sparsify activations but may introduce approximation errors or hardware overhead. Alternatively, bio-inspired methods have explored efficient replacements for BP, such as forward-forward methods (Dellaferrera & Kreiman, 2022; Hinton, 2022) and local learning rules (Lillicrap et al., 2014; Nøkland, 2016; Apolinario et al., 2025a; Frenkel et al., 2021; Apolinario et al., 2025b). They usually incur accuracy degradation. More recently, low-rank approaches exploit redundancy in activations using SVD or Tucker decompositions (Nguyen et al., 2024; 2025), but these typically recompute decompositions at each training step, adding significant overhead and making them unsuitable for continual learning, where repeated factorization across tasks would require storing large task-specific matrices.
@@
A key insight from orthogonal gradient projection methods in continual learning (Saha et al., 2021; Liang & Li, 2024; Apolinario et al., 2024) is that neural activations often lie in a stable, low-rank subspace that can be reused across training. This suggests a natural question: Can we identify a compact activation subspace and reuse it throughout fine-tuning, while simultaneously enabling efficient continual learning?
@@
With the above in mind, we introduce LANCE (Low-rank Activation Compression), a framework that applies one-time higher-order singular value decomposition (HOSVD) to extract a reusable low-rank activation subspace, as shown in Fig. 1b. Unlike prior iterative approaches (Nguyen et al., 2024; 2025), LANCE computes the decomposition only once at the beginning of training. Subsequent activations are projected onto this fixed subspace throughout training, avoiding repeated decompositions, thereby, significantly reducing both memory usage, computational cost, and hardware overhead. In contrast to system-level methods such as TinyTL (Cai et al., 2020) or the 256KB training engine (Lin et al., 2022a), which freeze layers or restrict updates to fit memory budgets, LANCE enables full backpropagation across all layers by directly compressing stored activations. Importantly, these fixed subspaces also allow assigning new tasks to orthogonal components, inherently supporting on-device continual learning without storing large task-specific matrices as in previous continual learning (CL) methods (Saha et al., 2021; Saha & Roy, 2023; Apolinario et al., 2024; Lin et al., 2022b; c). By eliminating repeated decompositions while preserving full-model fine-tuning, LANCE provides a practical and scalable solution for memory- and energy-efficient on-device continual learning.
@@ 2.1 Preliminaries
Training deep neural networks requires storing intermediate activations during the forward pass so that they can be reused for gradient computation in backpropagation, as illustrated in Fig. 1a. Let $f({\mathbf{x}};{\bm{\theta}})$ denote a neural network with input $\mathbf{x}$} and parameters $\bm{\theta}$}. During forward propagation, each layer $l$ produces an activation tensor ${\bm{\mathsfit{X}}}^{(l)}$ that is stored in SRAM for later use in computing the gradient: $\nabla_{{\bm{\theta}}^{(l)}}\mathcal{L}\;=\;\frac{\partial\mathcal{L}}{\partial{\bm{\mathsfit{X}}}^{(l)}}\;\frac{\partial{\bm{\mathsfit{X}}}^{(l)}}{\partial{\bm{\theta}}^{(l)}}$ , where $\mathcal{L}$ is the loss function. The cumulative memory footprint of storing $\{{\bm{\mathsfit{X}}}^{(l)}\}_{l=1}^{L}$ across all $L$ layers often dominates SRAM usage, especially in convolutional networks. For example, on CIFAR-100 with ResNet18, storing activations during training can consume up to $5\times$ more memory than storing weights, quickly exhausting the limited resources of edge devices. A natural way to mitigate this bottleneck is to exploit the low-rank structure of activations. Instead of storing the full tensor ${\bm{\mathsfit{X}}}^{(l)}$, we can approximate it with a compressed representation and reconstruct (or use) it during backpropagation. While classical singular value decomposition (SVD) is defined for matrices, activations are inherently higher-order (e.g., batch $\times$ channels $\times$ height $\times$ width), making tensor decompositions more suitable.
@@ 2.2 Proposed Method: LANCE
4:  Estimate covariance ${\bm{B}}_{i}$ from calibration activations via (2)
@@
8:   Retain top $r_{d}$ columns of $\hat{{\bm{U}}}_{d}$ satisfying (5); set ${\bm{U}}_{d}\leftarrow\hat{{\bm{U}}}_{d}[:,1\!:\!r_{d}]$
@@
11:   Retain top $r_{i}$ columns of ${\bm{U}}_{i}$ satisfying (3)
@@
20: Estimate last-mode covariance ${\bm{B}}^{t}_{d}$ from $N_{\text{CL}}$ calibration mini-batches as in (2)
@@
The compressed ${\bm{\mathsfit{G}}}^{(l)}$ is saved instead of the full activation and reused during backpropagation. Unlike prior works that recompute low-rank decompositions every iteration, LANCE computes the projection matrices only once at the beginning of fine-tuning and reuses them across all epochs. This one-shot approach avoids expensive per-iteration SVDs and is particularly well-suited for energy-constrained devices. The method, summarized in Algorithm 1, has two phases: (1) a one-shot subspace calibration performed offline, and (2) on-device fine-tuning with low-rank activations, which projects activations into low-rank subspaces during the forward pass and computes gradients during the backward pass using the compressed activations (Fig. 1b). Each step is detailed below.
@@
During fine-tuning with BP (Fig. 1b), we distinguish the forward and backward passes.
Forward pass. Each activation ${\bm{\mathsfit{X}}}^{(l)}$ is projected into the learned low-rank subspace using the precomputed matrices $\{{\bm{U}}^{(l)}_{i}\}_{i=1}^{d}$, following (1) to obtain a core tensor ${\bm{\mathsfit{G}}}^{(l)}$. Only the compact tensor ${\bm{\mathsfit{G}}}^{(l)}$ and the fixed $\{{\bm{U}}^{(l)}_{i}\}_{i=1}^{d}$ need to be stored for backpropagation. This reduces the memory footprint from $\mathcal{O}(\prod_{i=1}^{d}n_{i})$ to $\mathcal{O}(\prod_{i=1}^{d}r_{i}+\sum_{i=1}^{d}n_{i}r_{i})$ per activation.
@@ 2.3 LANCE for On-Device Continual Learning
At the end of training on task $t$, we estimate the subspace ${\bm{M}}^{t}$ by accumulating the covariance of activations along the last mode, ${\bm{B}}_{d}^{t}$, as in (2) using $N_{\text{CL}}$ calibration mini-batches. An SVD decomposition of ${\bm{B}}^{t}_{d}$ yields a basis of principal directions, and we retain the top $r_{d}$ components that capture at least an energy fraction $\varepsilon_{\text{CL}}$. To ensure orthogonality with previously stored directions, we project ${\bm{B}}^{t}_{d}$ onto the complement of ${\bm{M}}^{t-1}$: $\hat{\bm{B}}^{t}_{d}\;=\;{\bm{B}}^{t}_{d}-{\bm{M}}^{t-1}{\bm{M}}^{t-1\top}{\bm{B}}^{t}_{d}$, where ${\bm{M}}^{0}$ is initialized as empty. We select the smallest $r_{d}$ such that:
@@
For a new task t+1t{+}1, we update the one-shot calibration of LANCE by incorporating the memory constraint only in mode-$d$; all other modes follow the procedure in Section 2.2. We compute the covariance ${\bm{B}}_{d}$ for the new task, remove contributions along ${\bm{M}}^{t}$, $\hat{\bm{B}}_{d}\;=\;{\bm{B}}_{d}-{\bm{B}}_{d}{\bm{M}}^{t}{\bm{M}}^{t\top}$, and select the top $r_{d}$ eigenvectors of $\hat{\bm{B}}_{d}$ that satisfy an energy threshold $\varepsilon$ similar to (5). The resulting low-rank factors used during fine-tuning are $\{{\bm{U}}^{(l)}_{i}\}_{i=1}^{d-1}\cup\hat{\bm{U}}^{(l)}_{d,[:,1:r_{d}]}$, where $\hat{\bm{U}}^{(l)}_{d,[:,1:r_{d}]}$ are the retained singular vectors of $\hat{\bm{B}}_{d}$. This guarantees that activations of the new task are compressed into a subspace orthogonal to those of previous tasks, thereby reducing catastrophic forgetting.
@@ Discussion.
Both memory update and null-space constrained calibration are performed offline, so they do not affect the efficiency of on-device fine-tuning. The method requires storing only the compact memory matrices $\{{\bm{M}}^{(l)}\}$ and applying inexpensive projections during calibration. By construction, interference with past tasks is eliminated along the stored subspaces, and forgetting is mitigated without replay buffers or large task-specific models. In contrast to prior gradient-projection methods (Saha et al., 2021; Saha & Roy, 2023; Lin et al., 2022b; c; Apolinario et al., 2024), which explicitly use $\{{\bm{M}}^{(l)}\}$ for online gradient projection and must store such matrices in SRAM during training, our approach achieves continual learning naturally by projecting activations into fixed low-rank subspaces that lie in the null space of previous tasks, retaining LANCE’s memory efficiency.
@@ 2.4 Convergence Analysis
Assumption A1 (Orthogonal truncation).
@@
Assumption A3 (Layer linearity in inputs).
@@
Assumption A4 (Energy capture).
One-shot HOSVD retains a fraction $\varepsilon\in(0,1]$ of activation energy per mode, inducing a bound on gradient leakage outside the retained input subspace (used in Theorem 2.2).
Theorem 2.1 (Projected gradient & descent).
Under A1 and A3, for any layer $l$ with input projector $\mathsf{P}^{(l)}$, the LANCE weight gradient equals the right-Frobenius orthogonal projection of the full gradient: $\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}=\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{full}}\,\mathsf{P}^{(l)}.$ Consequently, $\langle\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}},\,\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{full}}\rangle=\|\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}\|_{F}^{2}\geq 0$, so $\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}$ is a descent direction unless it vanishes.
Theorem 2.2 (Monotone decrease & projected stationarity).
Let $\mathsf{Q}$ be the block projector that right-multiplies each layer’s weight gradient by its $\mathsf{P}^{(l)}$. Under A2, the LANCE update ${\bm{\theta}}_{k+1}={\bm{\theta}}_{k}-\eta\,\mathsf{Q}\,\nabla\mathcal{L}({\bm{\theta}}_{k})$ with $\eta\in(0,1/L]$ satisfies $\mathcal{L}({\bm{\theta}}_{k+1})\;\leq\;\mathcal{L}({\bm{\theta}}_{k})\;-\;\tfrac{\eta}{2}\,\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}^{2},$ hence $\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}\to 0$ and every limit point $\bm{\theta}$$_{\star}$ obeys $\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{\star})=\mathbf{0}$.
{proposition}
[Stationarity gap vs. truncation] If along the iterates $\|(I-\mathsf{Q})\nabla\mathcal{L}({\bm{\theta}})\|_{2}\leq C\sqrt{1-\varepsilon}$ for some $C>0$, then any limit point $\bm{\theta}$$_{\star}$ satisfies $\|\nabla\mathcal{L}({\bm{\theta}}_{\star})\|_{2}\leq C\sqrt{1-\varepsilon}$.
@@ 3 Experimental Evaluation
(i) Datasets: For single-task fine-tuning, we use CIFAR-10/100, Oxford-IIIT Pets, Flowers-102, and CUB-200, covering a range of dataset sizes and complexities. For continual learning, we evaluate on Split CIFAR-100 (20 tasks), Split MiniImageNet (20 tasks), and the 5-Datasets benchmark (CIFAR-10, MNIST, SVHN, Fashion-MNIST, and notMNIST), following common CL protocols. (ii) Metrics: We report: classification accuracy on target datasets, memory usage measured as peak activation storage in MB (as estimated in Appendix C), computational cost in FLOPs, and for CL benchmarks, average accuracy (ACC) and backward transfer (BWT) as defined in (Apolinario et al., 2024).

### v2

@@ 1 Introduction
Several approaches have been proposed to address this issue. Checkpointing [chen2016checkpoint] reduces memory usage at the cost of extra recomputation; reversible networks [gomez2017reversible] eliminate the need to store activations but require specialized architectures; and quantization or activation pruning [chen2021actnn, barley2023activation, yu2022backrazor] compress or sparsify activations but may introduce approximation errors or hardware overhead. Alternatively, bio-inspired methods have explored efficient replacements for BP, such as forward-forward methods [dellaferrera2022pepita, hinton2022forward] and local learning rules [lillicrap2014random, nokland2016DFA, apolinario2025lls, Frenkel2021LearningNetworks, apolinario2025tess]. However, they usually incur accuracy degradation. More recently, low-rank approaches exploit redundancy in activations using SVD or Tucker decompositions [nguyen2024activation, nguyen2025beyond], but these typically recompute decompositions at each training step, adding significant overhead and making them unsuitable for continual learning, where repeated factorization across tasks would require storing large task-specific matrices.
@@
A key insight from orthogonal gradient projection methods in continual learning [Saha2021GradientLearning, liang2024inflora, apolinario2024code] is that neural activations often lie in a stable, low-rank subspace that can be reused across training. This suggests a natural question: Can we identify a compact activation subspace and reuse it throughout fine-tuning, while simultaneously enabling efficient continual learning?
@@
With the above in mind, we introduce LANCE (Low-rank Activation Compression), a framework that applies one-time higher-order singular value decomposition (HOSVD) to extract a reusable low-rank activation subspace, as shown in Figure˜1b. Unlike prior iterative approaches [nguyen2024activation, nguyen2025beyond], LANCE computes the decomposition only once at the beginning of training. Subsequent activations are projected onto this fixed subspace throughout training, avoiding repeated decompositions, thereby significantly reducing memory usage, computational cost, and hardware overhead. In contrast to system-level methods such as TinyTL [lin2020tinytl] or the 256KB training engine [lin2022ondevice], which freeze layers or restrict updates to fit memory budgets, LANCE enables full backpropagation across all layers by directly compressing stored activations. Importantly, these fixed subspaces also allow assigning new tasks to orthogonal components, inherently supporting on-device continual learning without storing large task-specific matrices as in previous continual learning (CL) methods [Saha2021GradientLearning, Saha2023ContinualProjection, apolinario2024code, Lin2022TRGP:Learning, Lin22CUBER, SD2023, lmsp2025]. By eliminating repeated decompositions while preserving full-model fine-tuning, LANCE provides a practical and scalable solution for memory- and energy-efficient on-device continual learning.
@@ 2.1 Preliminaries
Training deep neural networks requires storing intermediate activations during the forward pass so that they can be reused for gradient computation in backpropagation, as illustrated in Figure˜1a. Let $f({\mathbf{x}};{\bm{\theta}})$ denote a neural network with input $\mathbf{x}$} and parameters $\bm{\theta}$}. During forward propagation, each layer $l$ produces an activation tensor ${\bm{\mathsfit{X}}}^{(l)}$ that is stored in SRAM for later use in computing the gradient: $\nabla_{{\bm{\theta}}^{(l)}}\mathcal{L}\;=\;\frac{\partial\mathcal{L}}{\partial{\bm{\mathsfit{X}}}^{(l)}}\;\frac{\partial{\bm{\mathsfit{X}}}^{(l)}}{\partial{\bm{\theta}}^{(l)}}$ , where $\mathcal{L}$ is the loss function. The cumulative memory footprint of storing $\{{\bm{\mathsfit{X}}}^{(l)}\}_{l=1}^{L}$ across all $L$ layers often dominates SRAM usage, especially in convolutional networks. For example, on CIFAR-100 with ResNet18, storing activations during training can consume up to $5\times$ more memory than storing weights, quickly exhausting the limited resources of edge devices. A natural way to mitigate this bottleneck is to exploit the low-rank structure of activations. Instead of storing the full tensor ${\bm{\mathsfit{X}}}^{(l)}$, we can approximate it with a compressed representation and reconstruct (or use) it during backpropagation. While classical singular value decomposition (SVD) is defined for matrices, activations are inherently higher-order (e.g., batch $\times$ channels $\times$ height $\times$ width), making tensor decompositions more suitable.
@@ 2.2 Proposed Method: LANCE
4:  Estimate covariance ${\bm{B}}_{i}$ from calibration activations via Equation˜2
@@
8:   Retain top $r_{d}$ columns of $\hat{{\bm{U}}}_{d}$ satisfying Equation˜5; set ${\bm{U}}_{d}\leftarrow\hat{{\bm{U}}}_{d}[:,1\!:\!r_{d}]$
@@
11:   Retain top $r_{i}$ columns of ${\bm{U}}_{i}$ satisfying Equation˜3
@@
20: Estimate last-mode covariance ${\bm{B}}^{t}_{d}$ from $N_{\text{CL}}$ calibration mini-batches as in Equation˜2
@@
The compressed ${\bm{\mathsfit{G}}}^{(l)}$ is saved instead of the full activation and reused during backpropagation. Unlike prior works that recompute low-rank decompositions every iteration, LANCE computes the projection matrices only once at the beginning of fine-tuning and reuses them across all epochs. This one-shot approach avoids expensive per-iteration SVDs and is particularly well-suited for energy-constrained devices. The method, summarized in Algorithm 1, has two phases: (1) a one-shot subspace calibration performed offline, and (2) on-device fine-tuning with low-rank activations, which projects activations into low-rank subspaces during the forward pass and computes gradients during the backward pass using the compressed activations (Figure˜1b). Each step is detailed below.
@@
During fine-tuning with BP (Figure˜1b), we distinguish the forward and backward passes.
Forward pass. Each activation ${\bm{\mathsfit{X}}}^{(l)}$ is projected into the learned low-rank subspace using the precomputed matrices $\{{\bm{U}}^{(l)}_{i}\}_{i=1}^{d}$, following Equation˜1 to obtain a core tensor ${\bm{\mathsfit{G}}}^{(l)}$. Only the compact tensor ${\bm{\mathsfit{G}}}^{(l)}$ and the fixed $\{{\bm{U}}^{(l)}_{i}\}_{i=1}^{d}$ need to be stored for backpropagation. This reduces the memory footprint from $\mathcal{O}(\prod_{i=1}^{d}n_{i})$ to $\mathcal{O}(\prod_{i=1}^{d}r_{i}+\sum_{i=1}^{d}n_{i}r_{i})$ per activation.
@@ 2.3 LANCE for On-Device Continual Learning
At the end of training on task $t$, we estimate the subspace ${\bm{M}}^{t}$ by accumulating the covariance of activations along the last mode, ${\bm{B}}_{d}^{t}$, as in Equation˜2 using $N_{\text{CL}}$ calibration mini-batches. An SVD decomposition of ${\bm{B}}^{t}_{d}$ yields a basis of principal directions, and we retain the top $r_{d}$ components that capture at least an energy fraction $\varepsilon_{\text{CL}}$. To ensure orthogonality with previously stored directions, we project ${\bm{B}}^{t}_{d}$ onto the complement of ${\bm{M}}^{t-1}$: $\hat{\bm{B}}^{t}_{d}\;=\;{\bm{B}}^{t}_{d}-{\bm{M}}^{t-1}{\bm{M}}^{t-1\top}{\bm{B}}^{t}_{d}$, where ${\bm{M}}^{0}$ is initialized as empty. We select the smallest $r_{d}$ such that:
@@
For a new task t+1t{+}1, we update the one-shot calibration of LANCE by incorporating the memory constraint only in mode-$d$; all other modes follow the procedure in Section 2.2. We compute the covariance ${\bm{B}}_{d}$ for the new task, remove contributions along ${\bm{M}}^{t}$, $\hat{\bm{B}}_{d}\;=\;{\bm{B}}_{d}-{\bm{B}}_{d}{\bm{M}}^{t}{\bm{M}}^{t\top}$, and select the top $r_{d}$ eigenvectors of $\hat{\bm{B}}_{d}$ that satisfy an energy threshold $\varepsilon$ similar to Equation˜5. The resulting low-rank factors used during fine-tuning are $\{{\bm{U}}^{(l)}_{i}\}_{i=1}^{d-1}\cup\hat{\bm{U}}^{(l)}_{d,[:,1:r_{d}]}$, where $\hat{\bm{U}}^{(l)}_{d,[:,1:r_{d}]}$ are the retained singular vectors of $\hat{\bm{B}}_{d}$. This guarantees that activations of the new task are compressed into a subspace orthogonal to those of previous tasks, thereby reducing catastrophic forgetting.
@@ Discussion.
Both memory update and null-space constrained calibration are performed offline, so they do not affect the efficiency of on-device fine-tuning. The method requires storing only the compact memory matrices $\{{\bm{M}}^{(l)}\}$ and applying inexpensive projections during calibration. By construction, interference with past tasks is eliminated along the stored subspaces, and forgetting is mitigated without replay buffers or large task-specific models. In contrast to prior gradient-projection methods [Saha2021GradientLearning, Saha2023ContinualProjection, Lin2022TRGP:Learning, Lin22CUBER, apolinario2024code], which explicitly use $\{{\bm{M}}^{(l)}\}$ for online gradient projection and must store such matrices in SRAM during training, our approach achieves continual learning naturally by projecting activations into fixed low-rank subspaces that lie in the null space of previous tasks, retaining LANCE’s memory efficiency.
@@ 2.4 Convergence Analysis
Assumption A1 (Orthogonal truncation)
@@
Assumption A3 (Layer linearity in inputs)
@@
Assumption A4 (Energy capture)
One-shot HOSVD retains a fraction $\varepsilon\in(0,1]$ of activation energy per mode, inducing a bound on gradient leakage outside the retained input subspace (used in Proposition˜1).
Theorem 2.1(Projected gradient & descent)
Under Assumptions A1 and A3, for any layer $l$ with input projector $\mathsf{P}^{(l)}$, the LANCE weight gradient equals the right-Frobenius orthogonal projection of the full gradient: $\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}=\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{full}}\,\mathsf{P}^{(l)}.$ Consequently, $\langle\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}},\,\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{full}}\rangle=\|\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}\|_{F}^{2}\geq 0$, so $\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}$ is a descent direction unless it vanishes.
Theorem 2.2(Monotone decrease & projected stationarity)
Let $\mathsf{Q}$ be the block projector that right-multiplies each layer’s weight gradient by its $\mathsf{P}^{(l)}$. Under Assumption A2, the LANCE update ${\bm{\theta}}_{k+1}={\bm{\theta}}_{k}-\eta\,\mathsf{Q}\,\nabla\mathcal{L}({\bm{\theta}}_{k})$ with $\eta\in(0,1/L]$ satisfies $\mathcal{L}({\bm{\theta}}_{k+1})\;\leq\;\mathcal{L}({\bm{\theta}}_{k})\;-\;\tfrac{\eta}{2}\,\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}^{2},$ hence $\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}\to 0$ and every limit point $\bm{\theta}$$_{\star}$ obeys $\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{\star})=\mathbf{0}$.
Proposition 1(Stationarity gap vs. truncation)
If along the iterates $\|(I-\mathsf{Q})\nabla\mathcal{L}({\bm{\theta}})\|_{2}\leq C\sqrt{1-\varepsilon}$ for some $C>0$, then any limit point $\bm{\theta}$$_{\star}$ satisfies $\|\nabla\mathcal{L}({\bm{\theta}}_{\star})\|_{2}\leq C\sqrt{1-\varepsilon}$.
@@ 3 Experimental Evaluation
(i) Datasets: For single-task fine-tuning, we use CIFAR-10/100 [Krizhevsky2009LearningImages], Oxford-IIIT Pets [petsdataset], Flowers-102 [flower102dataset], CUB-200 [cub200], and ImageNet [Krizhevsky2012ImageNetNetworks], covering a range of dataset sizes and complexities. For continual learning, we evaluate on Split CIFAR-100 (20 tasks), Split MiniImageNet (20 tasks), and the 5-Datasets benchmark (CIFAR-10, MNIST, SVHN, Fashion-MNIST, and notMNIST), following common CL protocols. (ii) Metrics: We report the following metrics: classification accuracy in the target datasets; memory usage, measured as peak activation storage in MB (estimated in Appendix C); and computational cost in FLOPs. For continual learning benchmarks, we additionally report the average final accuracy over all tasks (ACC) and the Backward Transfer (BWT), which quantifies the forgetting of previously learned tasks when new tasks are introduced, as defined in Appendix A.2.